OpenStudy (anonymous):

@sauravshakya Find the distance between (-2, 8) and (11, 2). Round to the nearest hundredth

OpenStudy (tkhunny):

\[\sqrt{(11+2)^{2}+(2-8)^{2}}\]Ringing any bells?

OpenStudy (anonymous):

can u do it ? its really confusing 4 me :(

OpenStudy (tkhunny):

I did, already. The arithmetic is left to you.

OpenStudy (anonymous):

i got 133?

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

For what? 11+2 = 13 ----- 13^2 = 169 2-8 = -6 ----- (-6)^2 = 36 I'm not really seeing a 133 in there. How did you find yours?

OpenStudy (anonymous):

oh i put -36

OpenStudy (tkhunny):

Negative * Negative = Positive, right? Squaring Things is ALWAYS Positive. Where does that leave us?